Data archive is back. There may be a little lag in getting data from the last few nights.
data archive down until tuesday 10/9
Folks,
They are installing new power outlets in our server room. The computers, including mnem, will be down until 10/9.
RCT working
The mirror support compressor’s pressure switch, which activates and deactivates the pump, burned out. It was replaced with a new switch, one with a bit more control on its activation pressure, which should reduce the wear on this part.
The problem with the mirror support compressor seems to be resolved, and the telescope should be working normally. I will check the focus and pointing again tonight, then resume the scheduler.
